OpenStudy (anonymous):

Help Please!! Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2.

OpenStudy (anonymous):

let f(1)+f(2)= t = 7 then f(5)= t+f(4) f(4)= t+f(3) f(3)= t+(f(2)) just put the values

OpenStudy (anonymous):

Umm... so it 4?

OpenStudy (anonymous):

it's*

OpenStudy (anonymous):

talking about answer?

OpenStudy (anonymous):

Yes. Would the answer be f(5)=4n?

OpenStudy (anonymous):

nope 26

OpenStudy (anonymous):

Oh.. Okay